The ratio of boys to girls in a class is 7:5. There are 36 students in the class. How many students are girls?

Answer by ikleyn(52835)

.
In all, there are 7+5 = 12 equal groups of students in the class, that in total make 36 people.


So there are 3 students in each small group.


Answer.  There are 7*3 = 21 boys and 5*3 = 15 girls.

Solved.